OpenStudy (anonymous):

The concession stand sells hot dogs and drinks during their high school foot ball game. John bought 6 hot dogs and 4 drinks and paid $6.70. Jessica bought 4 hot dogs and 3 sodas and paid $4.65. What is the price of a drink? $0.55 $0.65 $0.75 $0.85

OpenStudy (anonymous):

What you want to do here since you have to unknowns (the price of the hot dog and the price of a drink), is create 2 equations. For example for John 6h + 4d = 6.70. Then create an equation for Jessica also. Then with one of the equations get one variable all alone on the left side of the equal sign, say the h, and get everything else on the right side of the equation, even the number in front of the h should be on the right, you want the h all by itself on the left. This equation will have a numerator with the price in it minus the number of drinks, and a denominator that was the number that was in front of the h. Now take that fraction, (all numbers and variables to the right of the equal sign), and substitute it for the h in the first formula and reduce. This will solve for d, then plug d into one of the original equations to solve for h. Your answers should be a hot dog = 75¢ and a drink = 55¢.

OpenStudy (angela210793):

x= the price of a hot dog and y=the price of a soda we form a system { 6x+4y=6.70 {4x+3y=4.65

OpenStudy (anonymous):

thanks darlings (:

OpenStudy (anonymous):

Your welcome :D